Can you help with the below?

14.    Calculating the GDS and TDS Ratios: Calculate the gross debt service (GDS) and the total debt service (TDS) ratios for the following data.

Monthly mortgage payment = $2100
Property taxes = $200
Heating costs = $115
Other housing costs = $70
Personal loan payment = $150
Car loan payment = $200
Credit card payment = $150
Gross monthly household income = $7800

A: GDS Calculation =

A: GDS Percentage Value =

A: TDS Calculation =

A: TDS Percentage Value =
Answer #1

GDS is the percentage of your income needed to pay all of your monthly housing costs including principal,interest,taxes and heat.

GDS CALCULATION:-

GDS=Monthly mortgage payment+Property taxes+Heating cost +Other housing costs /Gross  monthly income

GDS =$2100+$200+$115+$70/7800=0.3186

GDS PERCENTAGE =31.86%(APPROX)

TDS CALCULATION:-

Tds ratio is the income needed to cover all your debts.

TDS = All debts/Gross monthly income

=$2100+$200+$115+$70+$150+$200+$150/$7800

=0.3827

TDS %=0.3827*100=38.27%